How do you make a table of values and graph y<-x^2+x? Everytime i graph it i get a graph that looks nothing like the one in the back of my book i cant figure out the nubers to put in for x in the table of values because my teacher said to find the vertex and use it for the middle of the table but it is not working for me
Answer by richwmiller(17219)

so what does the graph in the book look like?
What is the vertex that you found?
x=-b/2a=
-1/2*(-1)=
1/2
y=-x^2+x
y=-1/4+1/2=1/4
The vertex is (1/2,1/4)
If you got something else that will throw your graph off.